# Do Androids Know They’re Only Dreaming of Electric Sheep?

### v1

@@ Abstract
We design probes trained on the internal representations of a transformer language model that are predictive of its hallucinatory behavior on in-context generation tasks. To facilitate this detection, we create a span-annotated dataset of organic and synthetic hallucinations over several tasks. We find that probes trained on the force-decoded states of synthetic hallucinations are generally ecologically invalid in organic hallucination detection. Furthermore, hidden state information about hallucination appears to be task and distribution-dependent. Intrinsic and extrinsic hallucination saliency varies across layers, hidden state types, and tasks; notably, extrinsic hallucinations tend to be more salient in a transformer’s internal representations. Outperforming multiple contemporary baselines, we show that probing is a feasible and efficient alternative to language model hallucination evaluation when model states are available.
@@ 1 Introduction
Do language models know when they’re hallucinating? [^1] That is, do they know that their electric sheep are ungrounded? Detecting hallucinations in grounded generation is commonly framed as a textual entailment problem. Prior work has largely focused on creating secondary detection models trained on and applied to surface text Ji et al. (2023). However, these secondary models are often of comparable complexity to the underlying language model that generated the error and ignore the information that was already computed during generation. In this work, we explore the degree to which probes on a decoder-only transformer language model’s feed-forward and self-attention states can detect hallucinations in a variety of grounded generation tasks.
Figure 1: Prediction of hallucination in a generated response by probing the hidden states of a transformer during decoding. The true annotated span of hallucination within the response is boxed in white; rows represent probes trained to detect hallucination from layers within a transformer’s hidden states, with columns representing their prediction for each token in a generated response.
In particular, as the language model generates a response, we probe its encoded tokens to determine whether the model has just begun to hallucinate or will eventually be found to have hallucinated. It is plausible that these encodings are “aware” of whether generation is retrieving information that is grounded in the prompt versus filling in plausible information that is stored in the model’s weights. Related work has studied how transformers retrieve relevant memories as they encode text Meng et al. (2022, 2023) and suggests that feed-forward layers act as associative memory stores Geva et al. (2021). A probe for hallucination detection must consider whether subsequent generated text is entailed by information in the prompt, given the retrieved memories, or is merely one plausible extrapolation from the prompt (see also Monea et al., 2023).
We develop three probes of increasing complexity across three grounded generation tasks: abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. For each task, we collect hallucinations in two ways: (1) from sampled responses, where we generate outputs from a large language model (LLM) conditioned on the inputs, or (2) by editing reference inputs or outputs to create discrepancies. We refer to these as organic and synthetic data, respectively. In both cases, we have human annotators mark hallucinations in the outputs. Method (2) produces hallucination annotations at a higher rate, though we find that the utility of these synthetic examples is lower as they do not come from the test distribution.
@@
  • We propose three probe architectures for detecting hallucinations and demonstrate improvements over multiple contemporary baselines in hallucination detection.
@@ 2 Related Work
We center our study of hallucinations in the setting of in-context generation Lewis et al. (2020) where grounding knowledge sources are provided within the prompt itself. This is in contrast to settings where a language model is expected to generate a response entirely from its world knowledge (i.e., grounded in the model’s training data) Azaria and Mitchell (2023).
Though the term has often appeared in reference to factuality Gabriel et al. (2021), hallucinations are not necessarily factually incorrect. In their survey, Ji et al. (2023) classify hallucinations as intrinsic, where generated responses directly contradict the knowledge sources, or extrinsic, where generated responses are neither entailed nor contradicted by the sources. We follow this taxonomy. For example, given the following task prompt for abstractive summarization,
@@
A generated summary sentence of “Deckard can now purchase an animal for his wife” is faithful, “Deckard can purchase a goat in Iran” is an intrinsic hallucination, and “Sheep are usually cuter and fluffier than goats” is an extrinsic hallucination. While models may leverage latent knowledge to inform their responses, such as recognizing that a goat is indeed an animal, directly stating such knowledge in a response—i.e. only generating “Goats are animals” in the example above—is often considered an extrinsic hallucination as it is neither entailed nor contradicted by a given knowledge source. Thus, non-hallucinatory generation in this setting involves balancing the language model’s retrieval and manipulation of knowledge, both from the given context and from its own stored weights.
@@
Abstractive …dialed down that initial TL;DR: The National Weather
Summarization forecast to predict 13 to 19 Service forecast called
named storms, including six for 13 to 19 named
to nine hurricanes… storms including six
to 11 hurricanes.
KGDG …B: ok how can I get a Three consecutive strikes Response that uses A: Indeed, you have to
turkey I think that may are known as a "turkey." the information in get four consecutive
be hard to get though knowledge: strikes to get one.
Data-to-Text name[The Wrestlers], Restaurant Located by the river-
priceRange[high] Description: side, The Wrestlers is
a restaurant with high
prices, and near Raja
Indian Cuisine.
@@
Lexical metrics commonly used in the evaluation of NLG fail to correlate with human judgments of faithfulness Kryscinski et al. (2019). More useful are NLI approaches that either directly measure the degree of entailment between a generated response and a given source text Hanselowski et al. (2018); Kryscinski et al. (2020); Goyal and Durrett (2020), or else break down texts into single-sentence claims that are easily verifiable Min et al. (2023); Semnani et al. (2023); these have been successfully used for hallucination detection Laban et al. (2022); Bishop et al. (2023). Similarly, approaches that leverage question-answer models to test if claims made by a response can be supported by the source text, and vice versa Wang et al. (2020); Durmus et al. (2020); Scialom et al. (2021), try to more explicitly identify the grounding sources of given generations in this task. Under the intuition that a model is usually not confident when it is hallucinating, hallucinatory behavior can also be reflected in metrics that measure generation uncertainty or importance Malinin and Gales (2021); Guerreiro et al. (2023); Ren et al. (2023); Ngu et al. (2023); Madsen et al. (2023), relative token contributions Xu et al. (2023), and differences between the attention matrices and decoder stability Lee et al. (2018) during generation.
@@
In small transformer language models, it is possible to ascribe behaviors like in-context learning to specific layers and modules Elhage et al. (2021). This becomes harder in larger language models, but Meng et al. (2022) provide evidence that the feed-forward modules function as key-value stores of factual associations and that it is possible to modify weights in these modules to edit facts associated with entities. Our probes, to be successful, must determine whether the recalled facts were deployed in accordance with the rules of grounding (i.e., to paraphrase facts expressed in the prompt, rather than to replace them or to add details that are merely probable). Transformer internal states have also been shown to be predictive of the likelihood that a model will answer a question correctly in the setting of closed-book QA Mielke et al. (2022) and in the prediction of the truthfulness of a statement Azaria and Mitchell (2023), though these internal representations of truthfulness can disagree systematically with measures of model confidence Liu et al. (2023). These works measure hidden states in the absence of supporting evidence (i.e. no passages are retrieved for either task), making them effectively probes of knowledge stored in model weights and not the manipulation of novel or contingent information provided in a prompt. Monea et al. (2023) find that there are separate computational processes in transformer LLMs devoted to recall of associations from memory vs grounding data, and we show that we can detect whether associations from grounding data are used correctly across several distinct generation tasks. Slobodkin et al. (2023) train a probe to detect the (un)answerability of a question given a passage, which is similar in some respects to our task in Section 4; however, we detect token-level hallucination during answer generation, which could happen even in theoretically answerable questions.
@@
Table 2: Examples of changed knowledge statements for our datasets of synthetic hallucinations. Changes made to Conv-FEVER (top) and E2E (bottom) are shown in bold. The changed statements make the top response an intrinsic predicate error and the bottom response an extrinsic entity error (highlighted as in Table˜1).
3 Grounded Generation Tasks
We test hallucination probes for autoregressive grounded generation in three distinct tasks:  abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. We use the same high-level prompt structure across all tasks to organically generate model responses to tasks and to probe for hallucination. Each prompt contains an optional interaction history, a knowledge source, and an instruction resulting in the conditionally generated response. Task prompts and examples of organic hallucinations are shown in Table˜1.
@@ 3.1 Organic Hallucinations
We use llama-2-13b Touvron et al. (2023) as the response model to generate outputs for each task. To ensure that a non-negligible number of both hallucinations and grounded responses were present in the generations for each task, each model response is generated using top-k random sampling Fan et al. (2018) with $k=2$. The number of few-shot in-context learning examples for each task was chosen through a pilot exploration of several generation strategies, optimizing for the presence of a decent balance of hallucinations and grounded responses in all tasks, as determined via manual inspection by the authors. Full prompt templates are shown in Appendix˜C.
@@
The CNN / DailyMail (CDM) dataset Hermann et al. (2015) is a standard benchmark in the setting of abstractive summarization Lin and Ng (2019). It consists of a set of news articles paired with reference summaries in the form of human-written lists of bullet points that were originally published alongside the articles. For automatic summarization, we follow the GPT-2 paper Radford et al. (2019) and prompt our model to generate abstractive summaries of news articles by appending TL;DR: after the article content (a 0-shot prompt), [^2] In Internet posts that appear in LLM pretraining data, this string often marks the start of a summary immediately following the main content of the post. Prior work Radford et al. (2019) has found that appending this string after texts in prompts to autoregressive LLMs is capable of producing abstractive summaries of the former. taking the first three generated sentences as the summary.
@@
The E2E task Novikova et al. (2017) asks respondents to write restaurant descriptions that are congruent with a provided set of attributes. Here, we prompt our response model to perform this task with 10-shot in-context learning. [^3] As the original E2E task had omitted or missing information in up to 40% of its data, we use the cleaned version by Dušek et al. (2019). As even the cleaned version had instances of omitted or missing information, 10 in-context examples were manually selected and manually verified for groundedness.
@@ 3.2 Synthetic Hallucinations
Recent work Ma et al. (2023); Monea et al. (2023) has carefully created synthetic ungrounded datasets by hand to evaluate metrics used to measure unfaithfulness and models used to detect it. The BUMP dataset Ma et al. (2023) is a version of CDM in which reference summaries have been manually edited to be ungrounded. It includes both freestyle edits and systematic coverage of 4 specific types of faithfulness errors: (1) predicate errors, where the predicate in a generation is inconsistent with the source knowledge; (2) entity errors, where the subject or object of a predicate is inconsistent; (3) circumstance errors, or errors in the time, duration, or location of an event of the predicate; and (4) coreference errors. Types (1)–(3) may be intrinsic or extrinsic.
@@
We use BUMP in our CDM experiments. For our Conv-FEVER and E2E experiments, we similarly created faithfulness errors in a uniformly sampled subset of these datasets. However, rather than changing an example’s reference response, we changed its provided knowledge sentence so that the existing reference response would be unfaithful to it. This simulates the common case of hallucination where a large language model generates statements that may be true, but which are not grounded in the source documents as required. For Conv-FEVER, we prompted ChatGPT to modify a given example’s provided knowledge sentence to introduce intrinsic and extrinsic errors of types (1)–(4) above; we then manually judged these attempts and edited the unsuccessful ones. [^4] In hindsight, this workflow proved inefficient; significant human editing and validation were needed. In this way, we created equal numbers of examples with intrinsic and extrinsic predicate, entity, circumstance, and coreference errors, as well as freestyle errors.
@@
Similarly, for E2E, we modified a random subset of the provided grounding attributes. Specifically, for an example with $n$ attributes, we drew $k$ uniformly from [1,n1][1,n-1], uniformly selected one of the subsets of size $k$, and then flipped a fair coin to decide whether to remove or perturb—by randomly sampling an alternative attribute from the entire dataset—these $k$ attributes. Finally, we manually verified that the reference response was unfaithful with respect to the modified attributes; if not, we further edited the attributes. Examples of changed knowledge statements are shown in Table˜2; statistics of our synthetic datasets are shown in Table˜3.
Task N-o H-o N-s H-s
CDM 3399 23% 1778555Obtained via BUMP Ma et al. (2023). 50%
CF 924 64% 2846 50%
E2E 4975 57% 9950 50%
@@
Table 3: Dataset statistics. N-o is the number of annotated organic generations, and H-o is the percentage of organic generations with at least one hallucination. N-s and H-s refer to the same for synthetic responses. Note that grounded synthetic responses were not generated by a model but come from gold reference data.

### v2

@@ Abstract
We design probes trained on the internal representations of a transformer language model to predict its hallucinatory behavior on three grounded generation tasks. To train the probes, we annotate for span-level hallucination on both sampled (organic) and manually edited (synthetic) reference outputs. Our probes are narrowly trained and we find that they are sensitive to their training domain: they generalize poorly from one task to another or from synthetic to organic hallucinations. However, on in-domain data, they can reliably detect hallucinations at many transformer layers, achieving 95% of their peak performance as early as layer 4. Here, probing proves accurate for evaluating hallucination, outperforming several contemporary baselines and even surpassing an expert human annotator in response-level detection F1. Similarly, on span-level labeling, probes are on par or better than the expert annotator on two out of three generation tasks. Overall, we find that probing is a feasible and efficient alternative to language model hallucination evaluation when model states are available.
@@ 1 Introduction
Do language models know when they’re hallucinating? [^1] That is, do they know their electric sheep are ungrounded? Detecting hallucinations in grounded generation tasks (such as document summarization) is commonly framed as a textual entailment problem Ji et al. (2023), with prior work largely focused on creating or applying secondary detection models trained on and applied to surface text (see Falke et al. (2019); Huang et al. (2020); Kryscinski et al. (2020); Goyal and Durrett (2020); Mishra et al. (2021); and others). In this work, we explore the degree to which the generation model itself—a decoder-only transformer—already encodes the desired diagnostic information in its feed-forward and self-attention output states. We train span-level probes supervised at the token-level and response-level probes for hallucination on several grounded generation tasks and show that hallucination with respect to grounding data is detectable from an transformer’s hidden states, in line with similar results on probing ungrounded (i.e., closed-book QA) generation Mielke et al. (2022); Azaria and Mitchell (2023).
Figure 1: Prediction of hallucination in one generated response by probing the hidden states of a transformer during decoding. The true annotated span of hallucination within the response is boxed in white; rows represent probes trained to detect hallucination from different transformer layers, with columns representing their prediction for each token in a generated response.
As a language model generates a response, we apply our probes to its encoded tokens to determine whether the current token is likely to be a hallucination. Related work has argued that these encodings are full of meta-information about the model’s generation process. For example, salient information for generation can be localized to specific model layers and time steps Geva et al. (2021); Meng et al. (2022, 2023), and hidden states can encode the model’s confidence about the correctness of subsequent response generation Mielke et al. (2022). This suggests that a probe on model hidden states will have access to a rich set of features for detecting hallucination—features that would not be directly and efficiently available to a secondary entailment or classification model seeing only the input and generated output tokens.
We develop three probes of increasing complexity across three grounded generation tasks: abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. For each task, we collect hallucinations in two ways: (1) from ordinary sampled responses, where we generate outputs from a large language model (LLM) conditioned on the inputs, or (2) by editing reference inputs or outputs to create discrepancies. We refer to these as organic and synthetic data, respectively. In both cases, we have human annotators mark hallucinations in the outputs. Method (2) produces hallucination annotations at a higher rate, though we find that the utility of these synthetic examples is lower as they do not come from the test distribution.
@@
  • We propose three probe architectures for detecting hallucinations and demonstrate improvements over multiple contemporary baselines while also surpassing an expert human annotator on response-level detection F1. On two out of three tasks, the best probe achieves over 90% F1.
  • We show probe performance on span-level detection is competitive with an expert human annotator, performing on par or better on two out of three generation tasks in our dataset.
@@ 2 Related Work
We center our study of hallucinations in the setting of in-context generation Lewis et al. (2020) where grounding knowledge sources are provided within the prompt itself. This is in contrast to settings where a language model is expected to generate a response entirely from its parametric knowledge (i.e., grounded in the model’s training data, as in Azaria and Mitchell (2023)). Though the term has often appeared in reference to factuality Gabriel et al. (2021), hallucinations—as we use the term—are not necessarily factually incorrect, only ungrounded. We follow the taxonomy of Ji et al. (2023): in their survey, they classify hallucinations as either intrinsic, where generated responses directly contradict the knowledge sources, or extrinsic, where generated responses are neither entailed nor contradicted by the sources. We follow this taxonomy. For example, given the following task prompt for abstractive summarization,
@@
a generated summary sentence of “Deckard can now purchase an animal for his wife” is faithful, “Deckard can purchase a goat in Iran” is an intrinsic hallucination, and “A Nubian goat is a British domestic goat breed” is an extrinsic hallucination. While models may leverage latent knowledge to inform their responses, such as recognizing that a Nubian goat is indeed a British domestic goat breed, directly stating such knowledge in a response—as in the last sentence above—is considered an extrinsic hallucination when it is neither entailed nor contradicted by the knowledge sources. Thus, non-hallucinatory generation in this setting involves balancing the language model’s retrieval and manipulation of knowledge, both from the given prompt context and from its parameters.
@@
Abs. Summ. …dialed down that initial forcast to predict 13 to 19 named storms, including six to nine hurricanes… TL;DR: The National Weather Service forecast called for 13 to 19 named storms including six to 11 hurricanes.
KGDG …B: ok how can I get a turkey I think that may be hard to get though Three consecutive strikes are known as a “turkey.” Response that uses the information in knowledge: A: Indeed, you have to get four consecutive strikes to get one.
D2T name[The Wrestlers], priceRange[high] Restaurant Description: Located by the riverside, The Wrestlers is a restaurant with high prices, and near Raja Indian Cuisine.
@@
Lexical metrics commonly used in the evaluation of NLG fail to correlate with human judgments of faithfulness Kryscinski et al. (2019). More useful are natural language inference (NLI) approaches that either directly measure the degree of entailment between a generated response and a given source text Hanselowski et al. (2018); Kryscinski et al. (2020); Goyal and Durrett (2020), or else break down texts into single-sentence claims that are easily verifiable Min et al. (2023); Semnani et al. (2023). These NLI methods have been successfully used for hallucination detection Laban et al. (2022); Bishop et al. (2024). Similarly, approaches that leverage question-answer (QA) models to test if claims made by a response can be supported by the grounding text have also been applied to measure hallucination Wang et al. (2020); Durmus et al. (2020); Scialom et al. (2021). To the extent that models are less confident when they are hallucinating, hallucination may also be reflected in metrics that measure generation uncertainty Malinin and Gales (2021); Guerreiro et al. (2023); Ren et al. (2023); Ngu et al. (2024); Madsen et al. (2023), relative token contributions to the overall output probability Xu et al. (2023), or signatures in the attention matrices during generation Lee et al. (2018); Yuksekgonul et al. (2023). Similarly, conflicts between contextual knowledge provided in a prompt and parametric knowledge as learned during pre-training can contribute to this behavior Xu et al. (2024).
@@
In small transformer language models, it is possible to ascribe behaviors like in-context learning to specific layers and modules Elhage et al. (2021). This becomes harder in larger models, but Meng et al. (2022) provide evidence that the feed-forward modules function as key-value stores of factual associations and that it is possible to modify weights in these modules to edit facts associated with entities. Transformer internal states have also been shown to be predictive of the likelihood that a model will answer a question correctly in the setting of closed-book QA Mielke et al. (2022) and in the prediction of the truthfulness of a statement Azaria and Mitchell (2023). Interestingly, these internal representations of truthfulness can disagree systematically with measures of model confidence Liu et al. (2023). These works measure hidden states in the absence of supporting evidence (i.e. no passages are retrieved for either task), making them effectively probes of knowledge stored in model weights and not the manipulation of novel or contingent information provided in a prompt. In contrast, this work examines related questions specifically in the context of grounded generation (e.g., retrieval augmented generation Lewis et al. (2020)).
Concurrently to this work, Monea et al. (2023) find that there are separate computational processes in transformer LLMs devoted to the recall of associations from memory versus grounding data. Slobodkin et al. (2023) train a probe to detect the (un)answerability of a question given a passage, which is similar in some respects to our task in § 4; however, we detect span-level hallucination during answer generation, which could happen even in theoretically answerable questions. [^2] Reviewers asked to distinguish our work from that of Azaria and Mitchell (2023) and Kadavath et al. (2022). For a detailed discussion of the differences, see Appendix D.
3 Grounded Generation Tasks
@@
Table 2: Examples of changed knowledge statements for our datasets of synthetic hallucinations. Changes made to Conv-FEVER (top) and E2E (bottom) are shown in bold. The changed knowledge statements make the top response an intrinsic predicate error and the bottom response an extrinsic entity error (highlighted as in Table 1).
We test hallucination probes for autoregressive grounded generation in three distinct tasks:  abstractive summarization, knowledge-grounded dialogue generation, and data-to-text. We use the same high-level prompt structure across all tasks to organically generate model responses to tasks and to probe for hallucination. Each prompt contains an optional interaction history, a knowledge source, and an instruction resulting in the conditionally generated response. Task prompts and examples of organic hallucinations are shown in Table 1.
@@ 3.1 Organic Hallucinations
We use llama-2-13b Touvron et al. (2023) as the response model to generate outputs for each task. To ensure that a non-negligible number of both hallucinations and grounded responses were present in the generations, each model response is generated using top-k random sampling Fan et al. (2018) with $k=2$ and temperature $1$. The number of few-shot in-context learning examples for each task (see below) was chosen through a pilot exploration, optimizing for the presence of a balance of hallucinations and fully grounded responses, as determined via manual inspection by the authors. Full prompt templates are shown in Appendix C.
@@
The CNN / DailyMail (CDM) dataset Hermann et al. (2015) is a standard benchmark in the setting of abstractive summarization Lin and Ng (2019). It consists of a set of news articles paired with reference summaries in the form of human-written bullet points that were originally published alongside the articles. For automatic summarization, we follow Radford et al. (2019) and prompt our model to generate summaries of news articles by appending TL;DR: after the article content (a 0-shot prompt), [^3] In Internet posts that appear in LLM pretraining data, this string often marks the start of a summary immediately following the main content of the post. Prior work Radford et al. (2019) has found that appending this string after texts in prompts to autoregressive LLMs is capable of producing abstractive summaries of the former. taking the first three generated sentences as the summary.
@@
The E2E task Novikova et al. (2017) asks respondents to write restaurant descriptions that are congruent with a provided set of attributes. Here, we prompt our response model to perform this task with 9-shot in-context learning, with in-context examples manually validated for groundedness. [^4] As the original E2E task had omitted or missing information in up to 40% of its data, we use the cleaned version by Dušek et al. (2019). As even the cleaned version had instances of omitted or missing information, 9 in-context examples were manually selected and verified for groundedness.
@@ 3.2 Synthetic Hallucinations
Recent work Ma et al. (2023); Monea et al. (2023) has carefully created synthetic ungrounded datasets by hand to evaluate metrics used to measure unfaithfulness and models used to detect it. The BUMP dataset Ma et al. (2023) is a version of CDM in which reference summaries have been manually edited to be ungrounded. It includes both freestyle edits and systematic coverage of four specific types of errors: (1) predicate errors, where the predicate in a generation is inconsistent with the source knowledge; (2) entity errors, where the subject or object of a predicate is inconsistent; (3) circumstance errors, or errors in the time, duration, or location of an event of the predicate; and (4) coreference errors. Types (1)–(3) may be intrinsic or extrinsic, while (4) may be ambiguous; as such, we don’t consider coreference errors in our analysis of intrinsic or extrinsic hallucinations.
@@
We use BUMP in our CDM experiments. For our Conv-FEVER and E2E experiments, we similarly created errors in a uniformly sampled subset of these datasets. However, rather than changing an example’s reference response, we changed its provided knowledge sentence so that the existing reference response would be unfaithful to it. This simulates the common case of hallucination where a large language model generates statements that may be plausible but are not technically grounded.
For Conv-FEVER, we prompted ChatGPT to modify a given example’s provided knowledge sentence to introduce intrinsic and extrinsic errors of types (1)–(4) above; we then manually judged these attempts and edited the unsuccessful ones. [^5] In hindsight, this workflow proved inefficient; significant human editing and validation were needed. In this way, we created equal numbers of examples with intrinsic and extrinsic predicate, entity, circumstance, and coreference errors, as well as freestyle errors.
@@
For E2E, we modified a random subset of the provided grounding attributes. Specifically, for an example with $n$ attributes, we drew $k$ uniformly from $[1,n-1]$, uniformly selected one of the subsets of size $k$, and then flipped a fair coin to decide whether to remove or perturb—by randomly sampling an alternative attribute from the entire dataset—these $k$ attributes. Finally, we manually verified that the reference response was unfaithful with respect to the modified attributes; if not, we further edited the attributes. Examples of changed knowledge statements are shown in Table 2; statistics of our synthetic datasets are shown in Table 3.
Task N-o H-o $\kappa$-o N-s H-s
CDM 4684 28% 0.39 1778666Obtained via BUMP Ma et al. (2023). 50%
CF 924 70% 0.50 2846 50%
E2E 4998 54% 0.55 9950 50%
@@
Table 3: Dataset statistics. N-o is the number of annotated organic generations, H-o is the percentage of organic generations with at least one hallucination in the final adjudicated dataset, and $\kappa$-o is the response-level Fleiss’s $\kappa$ of individual annotations before adjudication. N-s and H-s refer to the same for synthetic responses. Note that grounded synthetic responses were not generated by a model but come from gold reference data.